OpenStudy (anonymous):

solve for x log(2)(x-2)-4=-log(2)(x+4)

OpenStudy (anonymous):

\[ x = -{\frac {\ln \left( 2 \right) -2}{\ln \left( 2 \right) }} \]

OpenStudy (anonymous):

maybe try \[\log_2(x-2)+\log_2(x+4)=4\] so \[\log_2((x-2)(x+4))=4\] and then \[(x-2)(x+4)=2^4\]

OpenStudy (anonymous):

\[x^2+2x-8=16\] \[x^2+2x=24\] \[(x+1)^2=25\] \[x+1=5\] \[x=4\]

OpenStudy (anonymous):

you may notice that i ignored \[x+1=-5\] \[x=-6\] because you cannot take the log of a negative number. also if completing the square is annoying you can write \[x^2+2x-24=8\] and factor as \[(x-4)(x+6)=0\] etc